OpenStudy (anonymous):

A square piece of cardboard 1 foot on each side is to be cut and folded into a box without the top(small squares will be cut off from each corner). Find the dimensions of the box with the greatest volume

OpenStudy (anonymous):

okay. suppose x feet are cut from each side. each of the cut outs will have an area of x^2. that means the height is x feet. let the length be L and the width be W L = 1-2x W = 1-2x volume is LWH V = x(1-2x)^2 to find the greatest volume, find out dV/dx and set it to 0

OpenStudy (anonymous):

the heigh (h) would be x??

OpenStudy (anonymous):

yes, try to visualize it. take a flat sheet of paper, cut out square shapes from each edge and try making it into a open box shape

OpenStudy (anonymous):

i get the derivative as being : 1-8x+12x^2

OpenStudy (anonymous):

okay! I havent done it, I can only tell you the method. I trust you can take it from there.

OpenStudy (anonymous):

so, i got the derivative as being dv/dx=1-8x+12x^2, then i got its roots as being 1/6 and 1/5. How do i find out what dimensions maximize the value?

OpenStudy (anonymous):

substitute each in the volume and see which one gives a bigger volume

OpenStudy (anonymous):

can you check if what i did is right, cause i got fractions as the critical points

OpenStudy (anonymous):

yes, you should get fractions. the side of the cardboard is 1 foot. so x is less than 1 foot, so it is logical that your answers should be less than 1.

OpenStudy (anonymous):

okay so i pliugged those citical values: 1/5 and 1/6 into the orginal volume formula: v=(1-2x)^2*x. For 1/5 i got .072. For 1/6 i got .074. Does this mean that 1/6 will gets me the greats volume, which also means that the dimension of the box should be 1+1/6 by 1+1/6 ???

OpenStudy (anonymous):

yes, then 1/6 is the correct value for x the dimensions of the cardboard are 1 foot by 1 foot. why would it be 1+1/6 by 1+1/6?

OpenStudy (anonymous):

Well, the question as me to find the dimensions of the box with the greatest volume?

OpenStudy (anonymous):

so the dimensions are L,W and H which are 1-2x,1-2x and x

OpenStudy (anonymous):

but where does the 1/6 come inot play, by that i mean what does it reperesent?

OpenStudy (anonymous):

okay. do exactly as I say. take a sheet of paper. tell me when you have done that.

OpenStudy (anonymous):

okay i ahve a sheet of paper

OpenStudy (anonymous):

okay, cut it so that it is a square sheet of paper if it is a rectangular sheet. if it is already a square sheet of paper, leave it as it is.

OpenStudy (anonymous):

also, join me here so that it is easier to explain. click on this link http://www.twiddla.com/573702